# CLIP-QDA: An Explainable Concept Bottleneck Model

### v1

@@ 4.1 Experimental setup
We evaluate our methods on ImageNet (Deng et al., 2009), PASCAL-Part (Donadello & Serafini, 2016), MIT Indoor Scenes dataset (Quattoni & Torralba, 2009), Monumai (Lamas et al., 2021) and Flowers102 (Nilsback & Zisserman, 2008). In addition to these well-established datasets, we introduce a custom dataset named Cats/Dogs/Cars dataset. To construct this dataset, we concatenated three widely recognized datasets, namely, the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). Subsequently, we filtered the dataset to exclusively contain images of white and black animals and cars. This curation resulted in six distinct subsets: “Black Cars”, “Black Dogs”, “Black Cats”, “White Cars”, “White Dogs”, “White Cats”. The primary objective of this dataset is to facilitate experiments under conditions of substantial data bias, such as classifying white cats when the training data has only encountered white dogs and black cats. In its final form, the dataset comprises 6,436 images.
@@
In addition to CLIP-QDA, we also test our method with linear regression as a classifier from CLIP scores, frequently named linear probe classifier. The objective here is to provide a comparison with existing CLIP-CBMs (Yang et al., 2023; Yan et al., 2023a; b).
@@ 4.2 Gaussian prior hypothesis
Figure 5: Multivariate QQplot for a Gaussian fit for the PASCAL-Part images that have the label “Aeroplane”.
Notably employing a less precise set of concepts can introduce disturbances, as evidenced in the observations. As indicated in Section 3.2.2, the ambiguity associated with certain concepts, such as “Multi-doored”, can lead to bimodal distributions (an aeroplane having one, multiple, or no doors). In Figure 6, we shows the histogram of the clip scores ${\bm{z}}$ of the images that have the class “aeroplane”. Compared to the histogram of less ambiguous cases (like in 3), we observe that the histogram presents anomalies, especially around the mean.
@@
Figure 7: Multivariate QQplot for a Gaussian fit for the PASCAL-Part images that have the label “Aeroplane”.
4.3 Classification accuracy
@@
In this section, we undertake a comparative evaluation of the performance of our CLIP-QDA classifier in contrast to other classifiers commonly employed in contemporary research. Precisely, we compare our method to the use of CLIP as a zero-shot classifier, by making predictions based on the class that produces the maximum CLIP score with the image embedding. We also use the linear probe, which consists of using a linear layer as the classifier of the CBM (see Figure 2). Our assessment is based on the computation of accuracy metrics across three diverse datasets: PASCAL-Part, MIT scenes, and Monumai. A comprehensive description of the procedure for concept set acquisition is provided in Appendix A.1. Furthermore, for contextualization, we include in the table the test accuracy of ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2021) trained in a supervised manner by using images as inputs.
Table 1: Test set accuracy.
Method PASCAL-Part MIT scenes Monumai
CLIP (zero shot) 0,81 0,63 0,52
CLIP + linear probe 0,91 0,77 0,81
CLIP + QDA probe 0,90 0,81 0,89
Resnet 50 (He et al., 2016) 0,84 0,86 0,95
ViT-L 336px (Dosovitskiy et al., 2021) 0,95 0,94 0,94
@@
Figure 8: Accuracy of the classifiers for different concept sizes. QDA refers to the use of CLIP-QDA. Linear probe refers to the use of a linear layer as a classifier.
4.4 XAI metrics
4.4.1 Global metric
To evaluate the relevance of our explanations, we present a toy example from our Cats/Dogs/Cars dataset by constructing a CBM consisting of the concepts Table 2, plus the concepts “Black” and “White”. Next, we display the 10 most influential concepts according to our global metric (the top 10 concepts that have the highest Wasserstein distance).
We present the results in Figure 9 for two scenarios: one with an unbiased dataset containing cats and cars of both colors and the other one with only black cats and white cars. We can observe that in the biased dataset, the concepts “Black” and “White” hold significantly higher importance, indicating that the classifier is likely to be biased about these concepts. This shows that our global explanation method has a potential for detecting biases in datasets (Tommasi et al., 2017). Additional explanation samples for various use cases are available in Appendix A.4.
(a) Unbiased dataset
(b) Biased dataset
@@ 4.4.1 Global metric
4.4.2 Local metrics
Test on a biased dataset.
In this section, we show an application of our local metric within the framework of our biased dataset, as previously described. We keep the same setup, employing a training set consisting of white cats and black cars. Subsequently, we feed an image of a white cat into our classifier (Figure 10). It is noteworthy that the image is misclassified as a car. Our local metric demonstrates sensitivity to the dataset’s color bias, corroborating the warning issued by the global explanation.
@@ 4.4.2 Local metrics
(b) Top 5 concepts. Classifier inference=“Car”
@@
Figure 10: Local explanation on subsets of Cats/Cars. On the right figure, the blue/red scale represents the scaled counterfactuals as in equation 6, the text in each box corresponds to the label predicted after the perturbation, followed to the concept changed to obtained its result (in parentheses). According to the explanation, we can note that removing a little of the concept “White”, or adding a little of the concept “Black” to the vector ${\bm{z}}$ induces a change of label to the correct one, thus highlighting the bias of the dataset.
Comparison to post-hoc methods.
We compare our local explanation method with popular post-hoc methods, namely LIME (Ribeiro et al., 2016) and SHAP (Lundberg & Lee, 2017). This involves treating CLIP scores as tabular data to generate explanations for the classifier’s decisions. These post-hoc methods can be more time-consuming, as illustrated in Figure A.3. This increased computational cost arises from their approach of analyzing the model’s behavior based on the output response to various perturbed inputs, rather than directly examining the model’s parameters.
To conduct this evaluation, we designed an experiment based on the deletion metric (Petsiuk et al., 2018). The procedure involves taking each sample from the test set and nullifying (i.e., setting to the average value of the score across classes) a certain number, $N_{null}$, of concepts. We nullify the concepts based on their importance order as determined by each explanation method. If nullifying the concepts leads to a significant decrease in performance, we consider it a successful selection of concepts that influenced the classifier’s decision. Our experiment is conducted using the unbiased version of the Cats/Dogs/Cars dataset. It is organized into two setups: the first one uses the concept set from Table 2, and the second one uses an equal number of concepts randomly selected from a dictionary of words. The results are presented in Figure 11.
(a) Random set of concepts
(b) Set of concepts that tailored to the dataset
Figure 11: Deletion test accuracies. The lower the accuracy is, the more faithful the explanation method is. Random refers to the use of a random selection of concepts to nullify.
In the random words setup, our local metric demonstrates competitive performance compared to LIME but lags behind SHAP. However, in the adapted word set scenario, we observed significantly worse results. This might be explained by the high correlation among the resulting CLIP scores from this word set, leading to situations akin to Figure 3(a), where searching for counterfactuals attribute by attribute struggled to provide meaningful insights. Globally, the classification boundaries obtained from our QDA are relevant for areas that present a high density of training samples, which is not the case for many of the counterfactuals.

### v2

@@ 4.1 Experimental Setup
We evaluate our methods on ImageNet (Deng et al., 2009), PASCAL-Part (Donadello & Serafini, 2016), MIT Indoor Scenes dataset (Quattoni & Torralba, 2009), MonuMAI (Lamas et al., 2021) and Flowers102 (Nilsback & Zisserman, 2008). In addition to these well-established datasets, we introduce a custom dataset named Cats/Dogs/Cars dataset (Section A.2). To construct this dataset, we concatenated two widely recognized datasets, namely, the Kaggle Cats and Dogs Dataset (Cukierski, 2013) and the Standford Cars Dataset (Krause et al., 2013). Subsequently, we filtered the dataset to contain images of white and black animals and cars exclusively. This curation resulted in six distinct subsets: “Black Cars”, “Black Dogs”, “Black Cats”, “White Cars”, “White Dogs”, “White Cats”. The primary objective of this dataset is to facilitate experiments under conditions of substantial data bias, such as classifying white cats when the training data has only encountered white dogs and black cats. Specifically, we refer to two distinct scenarios (Table 1): one containing cats and cars of both colors (referred to as the unbiased setup) and the other one with only black cats and white cars (referred to as the biased setup). When none of theses scenarios are referred (for example in the $Del$ values of Table 4 and 5), the complete dataset is used. In its final form, the dataset comprises 6,436 images. The dataset will be shared after the reviewing process.
Table 1: Cats/Dogs/Cars compositions used in our study
Complete dataset Biased setup Unbiased setup
Composition Black Cats, White Cats, Black Cars, White Cats, Black Dogs, White Dogs Black Cats, White Cars Black Cats, White Cats, Black Cars, White Cars
Num samples 6436 2536 4031
@@
In addition to CLIP-QDA, we also test our method with linear regression as a classifier from CLIP scores, frequently named linear probe classifier. The objective here is to provide a comparison with existing CLIP-based CBMs (Yang et al., 2023; Yan et al., 2023a; b).
4.2 CBM Quantitative Evaluation Process
Assessing the quality of explanations has long been a challenge, given its subjective nature. While quantitative evaluation processes exist for evaluating performance in a general setup (Hedström et al., 2023), they may not be well-suited to the unique characteristics of CBMs. Consequently, we present a novel method designed to evaluate XAI solutions tailored to CBMs. This method comprises two metrics: the Deletion metric, gauging faithfulness to the model, and the Detection metric, quantifying faithfulness to the data.
Deletion metric.
The deletion metric is adapted from the methodology introduced by Petsiuk et al. (2018). The procedure involves taking each sample from the test set and nullifying (i.e., setting to the average value of the score across classes) a certain number, $N_{null}$, of concepts. We nullify the concepts based on their importance order as determined by each explanation method. If nullifying the concepts leads to a significant decrease in performance, we consider it a successful selection of concepts that influenced the classifier’s decision. The intuition behind this idea is that if the concept is important, its absence will result in a loss of performance. Given the different values of accuracy obtained by nullifying $N_{null}$ concepts, we note $Acc(N_{null})$ we deduce the deletion score $Del$ by computing the area under the curve of $Acc(N_{null})$. The interest is to probe the ability to correctly order the important concept in the explanation:
$Del=\frac{1}{Acc(0)}\sum_{i=1}^{N_{max}}\frac{Acc(i-1)+Acc(i)}{2}\,.$ (7)
Here, $N_{max}$ represents a hyperparameter that defines the maximum number of deletions considered. The selection of this hyperparameter is critical: it balances between capturing the metric’s capacity to fit the model and ensuring that the computed inputs remain plausible. A default value of 9 is assigned to maintain this equilibrium. Note that we normalize the result by the maximum accuracy $Acc(0)$ to allow better comparisons among different classifiers.
The evaluation framework is structured into two setups: the first one uses the concept set outlined in Table 6 (referred to as Set 1), while the second one uses an equivalent number of concepts randomly chosen from a dictionary of words (referred to as Set 2).
Detection metric.
The detection metric evaluates the model’s ability to identify relevant concepts. For each sample $s$, a set of ground truth concepts $\mathcal{S}_{s}$ to detect is defined by an oracle. . Then, we construct a set of concepts $\mathcal{T}_{s}$ consisting of the top $|\mathcal{S}_{s}|$ concepts in the explanation to be tested. The resulting detection metric is the average ratio of agreements among all samples in the test dataset:
$Det=\frac{1}{S}\sum_{s=1}^{S}\frac{|\mathcal{S}_{s}\cap\mathcal{T}_{s}|}{|\mathcal{S}_{s}|},$ (8)
where $S$ is the total number of samples in the dataset. The selection of ground truth concepts depends on the available options within each dataset, as elaborated in Section 4.6.
@@ 4.3 Gaussian Prior Hypothesis
Figure 6: Multivariate Q-Q plot for a Gaussian fit for the PASCAL-Part images that have the label “Aeroplane”.
Notably employing a less precise set of concepts can introduce disturbances, as evidenced in the observations. As indicated in Section 3.2.2, the ambiguity associated with certain concepts, such as “Multi-doored”, can lead to bimodal distributions (an airplane having one, multiple, or no doors). In Figure 7, we shows the histogram of the clip scores ${\bm{z}}$ of the images that have the class “aeroplane”. Compared to the histogram of less ambiguous cases (like in 3), we observe that the histogram presents anomalies, especially around the mean.
@@
Figure 8: Multivariate Q-Q plot for a Gaussian fit for the PASCAL-Part images that have the label “Aeroplane”.
4.4 Assessing the Accuracy of XAI Methods
@@
In this section, we undertake a comparative evaluation of the performance of our CLIP-QDA classifier in contrast to other classifiers commonly employed in contemporary research. Precisely, we compare our method to the use of CLIP as a zero-shot classifier, by making predictions based on the class that produces the maximum CLIP score with the image embedding. We also use the Yan et al. (2023a) method, which consists of using a linear layer as the classifier of the CBM (see Figure 2). Additionally, we use Yang et al. (2023)’s method, involving the use of a class-concept matrix as the classifier of the CBM. Our assessment is based on the computation of accuracy metrics across three diverse datasets: PASCAL-Part, ImageNet, MIT scenes, and MonuMAI. A comprehensive description of the procedure for concept set acquisition is provided in Appendix A.3. All the CLIP-based CBMs are trained on the same sets of concepts denoted in Table 6. Furthermore, for contextualization, we include in the table the test accuracy of ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2021) trained in a supervised manner by using images as inputs. We also include Greybox XAI (Bennetot et al., 2022), X-NeSyL (Díaz-Rodríguez et al., 2022) and Morales Rodríguez et al. into our study, which offer explainability but require training data with annotated concepts. It is important to note that these methods, which require training from images, have significantly longer training times.
Results of these experiments are available in Table 2. The image encoder used in all our experiments based on CLIP-based CBMs is $ViT-L/14@336px$ provided from OpenAI’s public repository.
Table 2: Test set accuracy. On the top, are methods that require full training on images, and on the bottom, CBMs. Greybox XAI, X-NeSyL and Morales Rodríguez et al. requiring concept annotations, results on MIT scenes and ImageNet are not available.
Method PASCAL-Part    $\uparrow$ MIT scenes    $\uparrow$ MonuMAI    $\uparrow$ ImageNet    $\uparrow$
Non XAI Resnet 50 (He et al., 2016) 0.84 0.86 0.95 0.80
ViT-L 336px (Dosovitskiy et al., 2021) 0.95 0.94 0.94 0.85
CBMs Greybox XAI (Bennetot et al., 2022) 0.88 - 0.94 -
X-NeSyL (Díaz-Rodríguez et al., 2022) 0.82 - 0.90 -
Morales Rodríguez et al. 0.86 - 0.98 -
CLIP (zero-shot) (Radford et al., 2021) 0.81 0.63 0,52 0.76
LaBo (Yang et al., 2023) 0.83 0.75 0.74 0.69
Yan et al. (2023b), Yan et al. (2023a) 0.91 0.77 0.77 0.81
      CLIP-QDA (ours) 0.90 0.81 0.89 0.60
@@
Figure 9: Accuracy of the classifiers on PASCAL-Part for different concept sizes. QDA refers to the use of CLIP-QDA as a classifier. Linear probe refers to the use of a linear layer as a classifier.
4.5 Assessing the Dataset-wise Interpretability of XAI Methods
4.5.1 Evaluation of CLIP-QDAlocal and CLIP-QDAglobal on the Cats/Dogs/Cars dataset
To evaluate the relevance of our explanations, we present a toy example from our Cats/Dogs/Cars dataset by constructing a CBM consisting of the concepts Table 6, plus the concepts “Black” and “White”. Next, we display the 10 most influential concepts according to our global metric (the top 10 concepts that have the highest Wasserstein distance) in both biased and unbiased setups. Results are presented in Figure 10.
We can observe that the concepts “Black” and “White” hold significantly higher importance in the biased setup, indicating that the classifier is likely to be biased about these concepts. This shows that our global explanation method has the potential for detecting biases in datasets (Tommasi et al., 2017). Additional explanation samples for various use cases are available in Appendix A.7.
(a) Unbiased setup
(b) Biased setup
@@ 4.5.1 Evaluation of CLIP-QDAlocal and CLIP-QDAglobal on the Cats/Dogs/Cars dataset
We show an application of our local metric within the framework of our biased setup, as previously described. Subsequently, we feed an image of a white cat into our classifier (Figure 11). It is noteworthy that the image is misclassified as a car. Our local metric demonstrates sensitivity to the dataset’s color bias, corroborating the warning issued by the global explanation.
@@
(b) CLIP-QDAlocal explanation. Only the top 5 concepts are displayed. Note that the classifier misclassified the image as “Car”. The first row must be read as follows: removing a little of the concept “White” to the vector ${\bm{z}}$ induces a change of label to “Cat”.
@@
Figure 11: Local explanation on subsets of Cats/Cars. On the right figure, the blue/red scale represents the scaled counterfactuals as in equation 6, the text in each box corresponds to the label predicted after the perturbation, followed to the concept changed to obtain its result (in parentheses).
4.6 Assessing the Sample-wise Interpretability of XAI Methods
The objective of this subsection is to present our method within the context of current XAI methods applied to CBMs. We aim to compare CLIP-QDAglobal, CLIP-QDAlocal, CLIP-LIME and CLIP-SHAP with existing approaches such as SHAP (Lundberg & Lee, 2017), LIME (Ribeiro et al., 2016), LaBo (Yang et al., 2023), and Yan et al. (Yan et al., 2023b). A comparative overview of the distinctive features of these methods is presented in Table 3. Since Greybox XAI, X-NeSyL and Morales Rodríguez et al. require additional annotations, we chose to omit these methods from our study.
Table 3: Comparison of the different features of the XAI methods. Top: existing methods. Bottom: ours. Dataset-wise refers to methods that provide dataset-wise explanations. CLIP-QDAlocal refers to methods that provide sample-wise explanations. Closed-form solution refers to methods that do not require an optimization process to produce explanations. Computable from weights refers to methods that produce explanations from the model parameters and input values. Image level refers to methods that produce explanations from the image input. Concept level refers to methods that produce explanations from the concept input. Compatible with CLIP-QDA refers to methods that can produce explanations with CLIP-QDA as a classifier. Compatible with CLIP linear probe refers to methods that can produce explanations with CLIP linear probe as a classifier.
Metric Dataset wise Sample wise Closed-form solution Computable from weights Image level Concept level Compatible with CLIP-QDA Compatible with CLIP linear probe
GradCAM
LIME
SHAP
LaBo
Yan et al
Greybox XAI
X-NeSyL
Morales et al
CLIP-LIME
CLIP-SHAP
QDA-CBM
Qualitative analysis.
To facilitate the comparison of various explanations generated for the same sample in the dataset, we computed the results for one image from PASCAL-Part (Donadello & Serafini, 2016). The results are categorized into three sections: sample-wise explanations at the image level, sample-wise explanations at the concept level, and dataset-wise explanations at the concept level. The subsequent section focuses on explaining the model’s prediction process given the image 12(a) as the input, which is labeled as “person”. The top two predicted labels by the classifier for this image are “person” and “potted plant”. Note that additional samples are available in Section A.7.
Image-level explanations are presented in Figure 12. For each of these samples, a weight is assigned to each pixel. All the results are computed by applying these post-hoc explanation methods on CLIP-QDA. Additional implementation and computation details can be found in A.1. While the results are easy to comprehend, criticism may arise due to potential misunderstandings regarding the specific focus of the method. While these methods successfully highlight the object of interest, it is challenging to discern the exact pattern the method emphasizes. Another observation is that, compared to classical classifiers, the results on CLIP-based CBMs are more imprecise, highlighting large areas of interest.
(a) Input image
(b) GradCAM Explanation
(c) LIME Explanation
(d) SHAP explanation
Figure 12: Sample-wise explanations (image level). Note that the classifier classified the image correctly.
Concept-level sample explanations are calculated in Figure 13. On a sample-wise basis, CLIP-LIME and CLIP-SHAP are computed using CLIP-QDA. Given that the method is model-specific, the Yan et al. method is computed on a linear classifier.
(a) CLIP-QDAlocal explanation. The first row can be read as follows: removing a little of the concept “Human” to the vector ${\bm{z}}$ would change the label to “Car”.
(b) CLIP-LIME explanation. The first row can be read as follows: the fact that the concept score of “Facial features” is below 13.79 has a negative impact on the predicted label.
(c) Yan et al. explanation (sample). The first row can be read as follows: the concept “Posture or stance” has a positive impact on the predicted label.
(d) CLIP-SHAP explanation. The highest value can be read as follows: the concept “Fuselage” has a positive impact on the predicted label.
Figure 13: Sample-wise explanations (concept level) of the image 12(a). Except for CLIP-SHAP, only the top 5 concepts are displayed. Note that the classifier correctly labeled the image.
First, we notice that there is a high diversity of concepts displayed among the methods. Overall, there are concepts associated with the prediction label (“Posture or stance”, “Human”) and concepts that arise from the specificities of the image (“Leaves”). Some concepts seem irrelevant (“Landing Gear”, “Windshield”, “Wooly”), which may be explained by the fact that their scores are useful to the classification process despite their absence in the image. Plus, it is worth noting that the context provided by the counterfactual analysis helps to unveil insights into the reason behind the explanation. For example, when taking into account the concept “Leaves” seems inappropriate, CLIP-QDAlocal explanations suggest that adding that concept influences the prediction of the class “potted plant,” indicating that this concept is important because it reveals the decision process behind predicting a person rather than a plant.
In the realm of dataset-wise explanations (Figure 14), we present the explanations for the CLIP-QDAglobal method, computed on CLIP-QDA, the LaBo explanation, computed on the LaBo classifier, and the Yan et al. explanation, computed on a linear classifier.
(a) CLIP-QDAglobal explanation associated with the top two predictions (“person” and “potted plant”). The first row can be read as follows: the distribution of the concept “Container or pot” for the class “person” is mostly to the left (smaller values) of the one for the class “potted plant”.
(b) LaBo explanation associated with the top prediction (“person”). The first row can be read as follows: the weight associated with the concept “Displaying an image” has the highest value among the weights related to the class “person”.
(c) Yan et al. explanation (dataset) associated with the top prediction (“person”). The first row can be read as follows: the weight associated with the concept “Posture or stance” has the highest value among the weights related to the class “person”.
Figure 14: Dataset-wise explanations of the image 12(a). Only the top 5 concepts are displayed. Note that the classifier correctly labeled the image.
We observe that the concepts obtained from all the dataset-wise methods are generally coherent, including concepts commonly associated with persons, such as “Posture or stance” or “Facial features”. However, in contrast to the two other methods that impose positive magnitudes, the CLIP-QDAglobal explanation allows negative values, providing insights into whether the impact is negative or positive. Additionally, our CLIP-QDAglobal method emphasizes the comparison between two classes of interest, enabling a detailed exploration of specific disparities, like here between “potted plant” and “person”.
Quantitative analysis.
To conduct a more comprehensive analysis beyond qualitative observations, we subject the various methods to quantitative evaluations. We apply the procedure outlined in Section 4.2 to the PASCAL-Part and Cats/Dogs/Cars datasets. In addition to CLIP-LIME, CLIP-SHAP, and QDA-CBM explanations, we also tested the explanations provided by LaBo and Yan et al. For the selection of $S_{s}$, in the case of the PASCAL-Part dataset, we opted to designate the concepts to detect as the concepts that are associated with the two highest probabilities of the model’s inference (refer to the Classes/Concepts association in Table 6). In the Cats/Dogs/Cars dataset, we chose to identify the concepts "Black" and "White" in the biased setup to mitigate potential biases. The resulting $Det$ score for this dataset is calculated as the average across all possible biased binary classification tasks. Additionally, we incorporated inference time as a parameter in our experiments referring to the time taken to produce explanations inferences on the entire validation set. Results on PASCAL-Part are displayed in Table 5 and results on Cats/Dogs/Car are displayed in Table 4.
Table 4: Quantitative results for different XAI methods (Cats/Dogs/Cars dataset). Top: existing methods. Bottom: ours. Del refers to the deletion score (7) on a random set of concepts (Set 1) and the set of concepts defined in 6 (Set 2). Det refers to the detection score (8). The explanations of the three upper methods being direct, they present no computation time.
Method Del (Set 1) $\downarrow$ Del (Set 2) $\downarrow$ Det $\uparrow$ Inference time (im/s) $\downarrow$
Yan et al. 0.6253 0.7522 0.3183 /
LaBo 0.5971 0.6224 0.2140 /
Random 0.8278 0.7474 0.1171 /
CLIP-QDAlocal (ours) 0.7609 0.5820 0.2724 3.01
CLIP-LIME (ours) 0.5397 0.5646 0.4042 76.69
CLIP-SHAP (ours) 0.4821 0.3831 0.3696 256.76
Table 5: Quantitative results for different XAI methods (PASCAL-Part dataset). Top: existing methods. Bottom: ours. Del refers to the deletion score (7) on a random set of concepts (Set 1) and the usual set of concepts defined in 6 (Set 2). Det refers to the detection score (8). The explanations of the three upper methods being direct, they present no computation time.
Method Del (Set 1) $\downarrow$ Del (Set 2) $\downarrow$ Det $\uparrow$ Inference time (im/s)$\downarrow$
Yan et al. 0.6968 0.8824 0.4157 /
LaBo 0.4446 0.8822 0.4213 /
Random 0.8452 0.8999 0.1258 /
CLIP-QDAlocal (ours) 0.7983 0.8189 0.1564 2341.12
CLIP-LIME (ours) 0.7313 0.8511 0.0873 2857.28
CLIP-SHAP (ours) 0.5698 0.5510 0.2679 7207.58
Regarding the Cats/Dogs/Cars dataset, our methods exhibit superior performance compared to state of the art, emphasizing the reliability of CLIP-QDA on datasets with a low number of concepts. Particularly, QDA-CBM demonstrates significantly faster inference times than CLIP-LIME and CLIP-SHAP, albeit at the expense of slightly lower deletion and detection scores. On datasets with a higher number of concepts and classes, such as PASCAL-Part, existing methods maintain higher scores.